OpenStudy (anonymous):

856 seconds = how many min and seconds?

OpenStudy (lgbasallote):

divide 856 by 60...what do you get?

OpenStudy (anonymous):

14.266666666666666666666666667 ?

OpenStudy (helpmeplzz):

in this case, 856/60=14.2666666... so you have 14 whole minutes. 14x60=840 856-840=16

OpenStudy (lgbasallote):

so there's your number of minutes

OpenStudy (helpmeplzz):

Therefore you have a final answer of 14 minutes and 16 seconds.

OpenStudy (campbell_st):

divide by 60... you will get a mixed number so 14 x 60 = 840 seconds... your answer in minuteas and seconds will be 14 mins and (856 - 840= ) seconds..

OpenStudy (helpmeplzz):

@sara1234 so you have a final answer of 14 minutes and 16 seconds.